OpenStudy (clamin):

PLEASE HELP!! MEDAL!! Solve equation by factoring. 7x^2 + 13x =2 Solve equation by completing the square. r^2 - 6r - 80 = -8

OpenStudy (campbell_st):

rewrite the equation \[7x^2 + 13x - 2 = 0\] this can be factored, both 2 and 7 are prime so (7x - 1)(x + 2) = 0 now you can solve for x

OpenStudy (campbell_st):

for the 2nd equation, rewrite it as \[r^2 - 6r = 72\] now add \[(\frac{-6}{2})^2\] to both sides of the equation and you get \[r^2 - 6r + 9 = 81\] now factor the left side and you will be able to solve for r